OpenStudy (anonymous):

A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

will reward for help

OpenStudy (anonymous):

since 2x= 3-2y, this means 3x=4.5-3y. replace 3x in the first equation with 4.5-3y so we get 4.5-3y+8y=12, simplify to 5y=7.5, simplify again to y=1.5. plug this in to get x=0

OpenStudy (anonymous):

Thank u